Light passes from medium A into medium B.
As it does so the wavelength of the light is reduced by 22.0%.
What is the ratio of the refractive indices nB/nA? (to 3 s.f)
1
Expert's answer
2020-05-27T10:32:20-0400

According to Snell's Law, "\frac{sin \theta_1}{\sin \theta_2} = \frac{n_2}{n_1} = \frac{\lambda_1}{\lambda_2}", where n is a refractive index and "\lambda" is the wavelength. When passing from medium A into B, wavelength changes from "\lambda_1" to "(1-\eta)\lambda_1", therefore "\frac{n_2}{n_1}= \frac{\lambda_1}{(1-\eta)\lambda_1} = \frac{1}{1-\eta}". Substituting "\eta = 0.22", obtain "\frac{n_2}{n_1} \approx 1.282".